Aylward Inc. currently has ​$2,175,000 in current assets and ​$839,000 in current liabilities. The​ company's managers want to increase the​ firm's inventory, which will be financed by a​ short-term note with the bank. What level of inventories can the firm carry without its current ratio falling below 2.12.1​? The cost of the additional inventory financed with the​ short-term note is ​-----(Round to the nearest​ dollar.)
Answer #1

Answer

  • Purchasing Inventory through a short term note would increase Current Assets and Current Liabilities.
  • Let the amount of purchase be ‘x’
  • Required Current ratio = 2.12 to 1
  • Hence,

(2.12/1) = ($2175000 + x) / ($839000 + x)
2.12 x (839000 + x) = 2175000 + x
1778680 + 2.12x = 2175000 + x
2.12x – x = 2175000 – 1778680
1.12x = 396320
x = 396320 / 1.12
x = $ 353,857

  • Answer: The cost of the additional inventory financed with the​ short-term note is $ 353,857